OpenStudy (haleyelizabeth2017):

Solve for all radian solutions (showing algebra) 3 cos A + 3 = sin^2 A Please just walk me through the steps (:

OpenStudy (mathstudent55):

Hint: Use a trig identity that will give you cos^2 A. Then you will have a quadratic in cos A.

OpenStudy (alekos):

change sin^2(A) to 1 - cos^2(A)

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (alekos):

Then substitute x = cosA

OpenStudy (haleyelizabeth2017):

Huh?

OpenStudy (alekos):

and you'll have your quadratic

OpenStudy (alekos):

solve for x

OpenStudy (alekos):

A = arccos x

OpenStudy (alekos):

give it a try and see how you go

OpenStudy (haleyelizabeth2017):

So for the cos^2 A I'd do x^2?

OpenStudy (alekos):

yes

OpenStudy (haleyelizabeth2017):

Okay, so I got for the quadratic equation \[x^2+3x+2=0\] I know how to solve using the quadratic formula XD so just need to double check my equation

OpenStudy (alekos):

yes

OpenStudy (alekos):

just factorise

OpenStudy (haleyelizabeth2017):

I got -1 and -2.... @sleepyjess do you mind double checking? Thanks.

OpenStudy (mathstudent55):

That is x. Now substitute back cos A for x.

OpenStudy (alekos):

Yes, that's it and only one value is valid

OpenStudy (haleyelizabeth2017):

I'm confused...substitute back cos A for x? as in cos A = {-1, -2} ? Like I said, I'm confused LOL

OpenStudy (mathstudent55):

Remember that you used x for cos A? You got a quadratic in x and solved for x. x for you is really cos A. Now set cos A equal to the roots of the quadratic and solve the trig part.

OpenStudy (haleyelizabeth2017):

OH

OpenStudy (haleyelizabeth2017):

so arccos(-1) and arccos(-2)?

OpenStudy (haleyelizabeth2017):

I know arccos(-1) is 180 degrees, or pi radians...

OpenStudy (mathstudent55):

\(3 \cos A + 3 = \sin^2 A\) Use trig identity \(\sin^2 A + \cos^2 A = 1\) to get: \(3 \cos A + 3 = 1 - \cos^2 A\) \(cos^2 A + 3 \cos A + 2 = 0\) If you do not substitute and keep cos A, you have a quadratic in cos A, and you now factor: \((\cos A + 2)(\cos A + 1) = 0\) \(\cos A = -2\) or \(\cos A = -1\) \(\cos A = -2\) is outside of the range of the cosine function. The cosine function has a range in the interval [-1, 1]. That means the first root is discarded.

OpenStudy (mathstudent55):

You just need to solve \(\cos A = -1\) You are correct above.

OpenStudy (mathstudent55):

Remember you need all radian solutions, not just in the interval [0, 2pi)

OpenStudy (haleyelizabeth2017):

Okay, so how do we do that?

OpenStudy (mathstudent55):

You need to show that the solutions are \(\pi, \pi + 2\pi, \pi + 4\pi, \pi + 6\pi\), etc. and also \(\pi - 2 \pi\), \(\pi - 6 \pi\), etc.

OpenStudy (mathstudent55):

You use a variable to represent all integers, such as k or n, and you write the solution as \(A = \pi + 2\pi k\), for all integers k

OpenStudy (haleyelizabeth2017):

Okay, thank you!!!

OpenStudy (mathstudent55):

You're welcome.